# Dual Target-Mounted RISs-Assisted ISAC Against Eavesdropping and Malicious Interference

### v1

@@ II-C Communication Model
$\mathbf{H}_{n}=\sqrt{\frac{{L_{0}}}{d_{n}^{\alpha_{n}}}}\left({\sqrt{\frac{\kappa}{1+\kappa}}\mathbf{H}^{\text{LOS}}_{n}+\sqrt{\frac{1}{1+\kappa}}\mathbf{H}^{\text{NLOS}}_{n}}\right)$ (19)
$\mathbf{g}_{n,k}=\sqrt{\frac{{L_{0}}}{d_{n,k}^{\alpha_{k}}}}\left(\sqrt{\frac{\kappa}{1+\kappa}}\mathbf{g}_{n,k}^{\text{LOS}}+\sqrt{\frac{1}{1+\kappa}}\mathbf{g}_{n,k}^{\text{NLOS}}\right)$ (20)
@@
$\mathbf{H}^{\text{LOS}}_{n}=\mathbf{b}(\theta_{n}^{\text{RIS}},\phi_{n}^{\text{RIS}})^{\mathrm{H}}\mathbf{b}(\theta_{n}^{\text{BS}},\phi_{n}^{\text{BS}})$ (21)
$\mathbf{g}^{\text{LOS}}_{n,k}=\mathbf{b}(\theta_{n,k}^{\text{RIS}},\phi_{n,k}^{\text{RIS}})$ (22)
@@
$\mathbf{b}(\theta_{n}^{\text{BS}},\phi_{n}^{\text{BS}})=[1,\cdots,e^{j{\nu_{T}}(T_{x}-1)\cos(\theta^{\text{BS}}_{n})\cos(\phi^{\text{BS}}_{n})}]$ (23)
$\mathbf{b}(\theta_{n}^{\text{RIS}},\phi_{n}^{\text{RIS}})=\mathbf{b}^{x}(\theta_{n}^{\text{RIS}},\phi_{n}^{\text{RIS}})\otimes\mathbf{b}^{z}(\theta_{n}^{\text{RIS}},\phi_{n}^{\text{RIS}})$ (24)
where $\mathbf{b}^{x}(\theta_{n}^{\text{RIS}},\phi_{n}^{\text{RIS}})\in\mathbb{C}^{1\times N_{n}^{x}}$ and $\mathbf{b}^{z}(\theta_{n}^{\text{RIS}},\phi_{n}^{\text{RIS}})\in\mathbb{C}^{1\times N_{n}^{z}}$ represent the steering vectors of the $n$-RIS along the $x$- and $z$-axes, respectively. Here, $N_{x}^{n}$ and $N_{z}^{n}$ denote the number of reflecting elements along the $x$- and $z$-directions, and the total number of elements is given by $N_{n}=N_{x}^{n}\times N_{z}^{n}$ for $n\in\{L,M\}$. Therefore, $\mathbf{b}^{x}(\theta_{n}^{\text{RIS}},\phi_{n}^{\text{RIS}})$ and $\mathbf{b}^{z}(\theta_{n}^{\text{RIS}},\phi_{n}^{\text{RIS}})$ can be given as
$\mathbf{b}^{x}(\theta_{n}^{\text{RIS}},\phi_{n}^{\text{RIS}})=[1,\cdots,e^{j{\nu_{R}}(N_{n}^{x}-1)\cos(\theta^{\text{RIS}}_{n})\cos(\phi^{\text{RIS}}_{n})}]$ (25a)
$\mathbf{b}^{z}(\theta_{n}^{\text{RIS}},\phi_{n}^{\text{RIS}})=[1,\cdots,e^{j{\nu_{R}}(N_{n}^{z}-1)\sin(\phi^{\text{RIS}}_{n})}]$ (25b)
@@
$y_{k}(t)=\mathbf{g}_{L,k}\mathbf{\Omega}_{L}\mathbf{H}_{L}\mathbf{x}(t)+\mathbf{g}_{M,k}\mathbf{\Omega}_{M}\mathbf{H}_{M}\mathbf{x}(t)+n_{k}(t)$ (26)
@@
$\eta_{k}=\frac{\left\|\mathbf{g}_{L,k}\mathbf{\Omega}_{L}\mathbf{H}_{L}\mathbf{w}_{k}\right\|_{\mathrm{F}}^{2}}{\sum_{\hat{k}}\left\|\mathbf{g}_{L,k}\mathbf{\Omega}_{L}\mathbf{H}_{L}\mathbf{w}_{\hat{k}}\right\|_{\mathrm{F}}^{2}+\left\|\mathbf{g}_{M,k}\mathbf{\Omega}_{M}\mathbf{H}_{M}\mathbf{W}\right\|_{\mathrm{F}}^{2}+\sigma_{n}^{2}}$ (27)
@@ II-D Security Model
In the proposed scheme, since communication signals through users are assumed to be intercepted by an eavesdropper UAV, using (8), the received signal at the eavesdropper UAV target at the $t$-th time index can be expressed as
$y_{E}(t)=\sqrt{\beta_{E}}\mathbf{a}(\theta_{E}^{\text{BS}},\phi_{E}^{\text{BS}})^{\mathrm{H}}\mathbf{x}(t)+n_{E}(t)$ (28)
@@
$\eta_{E,k}=\frac{\left\|\mathbf{a}(\theta_{E}^{\text{BS}},\phi_{E}^{\text{BS}})^{\mathrm{H}}\mathbf{w}_{k}\right\|_{\mathrm{F}}^{2}}{\sum_{\hat{k}}\left\|\mathbf{a}(\theta_{E}^{\text{BS}},\phi_{E}^{\text{BS}})^{\mathrm{H}}\mathbf{w}_{\hat{k}}\right\|_{\mathrm{F}}^{2}+\sigma_{n}^{2}}$ (29)
@@
$S_{R,k}=[R_{k}-R_{E,k}]^{+}$ (30)
where $x]^{+}=\max(x,0)$. Here, $R_{k}=\log_{2}(1+\eta_{k})$ is the achievable rate of user Uk and $R_{E,k}=\log_{2}(1+\eta_{E,k})$ represents the eavesdropping rate at for Uk.
@@ III Problem Formulation and Proposed Solution
In the proposed scheme, the BS operates in dual mode for communication and sensing purposes, yet it is unaware of the malicious RIS mounted on the $E$-UAV. This malicious RIS forms its reflection matrix $\mathbf{\Omega}_{M}$ in (2) using randomly generated phase shifts. To counteract security threats of the $M$-RIS and its mounted $E$-UAV interception threats, the reflection coefficients of the legitimate RIS and the BS transmit beamforming matrix are optimized to maximize the system secrecy rate in (30). To address this, the following optimization problem is formulated:
$\text{(P1)}\hskip 11.38092pt\max_{\mathbf{W},\mathbf{\Omega}_{L}}\hskip 5.69046ptS_{R}$ (31a)
$\hskip 42.67912pt\text{s.t.}\hskip 5.69046pt\operatorname{Tr}(\mathbf{WW}^{\mathrm{H}})\leq P_{T}$ (31b)
$\hskip 65.44142pt\gamma_{i}\leq\rho_{s}\hskip 8.5359pt\text{for}\hskip 5.69046pti\in\{L,E\}$ (31c)
$\hskip 65.44142pt|e^{j\varphi_{l}}|=1,\hskip 8.5359pt\text{for}\hskip 5.69046ptl\in\{1,\cdots,N_{L}\}$ (31d)
where $S_{R}=\sum\limits_{k=1}^{K}S_{R,k}$ is the achievable sum secrecy rate of legitimate users and $\rho_{s}$ is the maximum achievable sensing SINR rate. Here, since both legitimate and eavesdropper UAVs communicate directly with the BS without any reflected signals from RISs, the beamforming optimization can simply be independent of $L$-RIS configuration. This allows the problem (P1) to be decomposed into a two-stage optimization framework that first optimizing the transmit beamforming matrix $\mathbf{W}$, and later optimizing the $L$-RIS reflection matrix $\mathbf{\Omega}_{L}$.
@@ III-A Beamforming Optimization
$\text{(P2)}\hskip 11.38092pt\max_{\mathbf{W}}\min\hskip 5.69046pt\gamma_{i}$ (32a)
$\hskip 56.9055pt{\text{s.t.}}\hskip 5.69046pt\gamma_{i}\leq\rho_{s}$ (32b)
$\hskip 76.82234pt\operatorname{Tr}(\mathbf{WW}^{\mathrm{H}})\leq P_{T}$ (32c)
Here, considering (8-9), for $\mathbf{A}_{i}\in\mathbb{C}^{T_{x}\times T_{x}}=\beta_{i}\mathbf{a}(\theta^{\text{BS}}_{i},\phi^{\text{BS}}_{i})^{\mathrm{H}}\mathbf{a}(\theta^{\text{BS}}_{i},\phi^{\text{BS}}_{i})$ being the round trip channel between BS and $i$-UAV target, the sensing SINR expression in (9) can be re-expressed as
$\gamma_{i}=\frac{\operatorname{Tr}(\mathbf{A}_{i}^{\mathrm{H}}\mathbf{A}_{i}\mathbf{R_{x}})}{\operatorname{Tr}(\mathbf{A}_{j}^{\mathrm{H}}\mathbf{A}_{j}\mathbf{R_{x}})+\sigma_{n}^{2}}$ (33)
where $j\neq i\in\{L,E\}$. Therefore, using a semi-definite relaxation (SDR)-based approach, (P2) can be reformulated as
$\text{(P2)}\hskip 11.38092pt\max_{\mathbf{R_{x}}}\hskip 2.84544pt\min\hskip 5.69046pt\gamma_{i}$ (34a)
$\hskip 54.06006pt{\text{s.t.}}\hskip 8.5359pt\gamma_{i}\leq\rho_{s}$ (34b)
$\hskip 73.97733pt\operatorname{Tr}(\mathbf{A}_{i}^{\mathrm{H}}\mathbf{A}_{i}\mathbf{R_{x}})\geq\operatorname{Tr}(\mathbf{A}_{j}^{\mathrm{H}}\mathbf{A}_{j}\mathbf{R_{x}})+\sigma_{n}^{2}$ (34c)
$\hskip 73.97733pt\operatorname{Tr}(\mathbf{R_{x}})\leq P_{T}$ (34d)
using the SeDuMi solver within the CVX optimization toolbox [29]. Subsequently, the beamforming matrix $\mathbf{W}$ can be directly obtained through eigenvalue decomposition (EVD) [17].
@@ III-B Legitimate RIS Phase Optimization
In this subsection, after the beamforming matrix $\mathbf{W}$ is obtained, the reflection coefficients of $L$-RIS are optimized to maximize the secrecy rate of the $K$ legitimate users. From (30), since the secrecy rate $S_{R}$ is the total sum of the achievable secrecy rates across all legitimate users, and eavesdropper SINR of $E$-UAV target (29) across users are independent from $L$-RIS reflection, the $\boldsymbol{\Omega}_{L}$ (1) can be optimized to enhance $S_{R}$ by maximizing sum of achievable rates of all users. Therefore, the optimization of $\mathbf{\Omega}_{L}$ in (P1) can be converted to following problem:
$\text{(P3)}\hskip 11.38092pt\max_{\mathbf{\Omega}_{L}}\hskip 2.84544pt\min\eta_{k}$ (35a)
@@
As a first step, the communication SINR given in (27) is re-expressed in a quadratic form by applying appropriate mathematical modifications. Here, applying trace equality [17], the numerator in (27) can be rewritten as $\left\|\mathbf{g}_{L,k}\mathbf{\Omega}_{L}\mathbf{H}_{L}\mathbf{w}_{k}\right\|_{\mathrm{F}}^{2}=\operatorname{Tr}(\mathbf{\bar{H}}_{L,k}\mathbf{\Omega}_{L}^{\mathrm{H}}\mathbf{\bar{G}}_{L,k}^{\mathrm{T}}\mathbf{\Omega}_{L})=\mathbf{z}_{L}(\mathbf{\bar{H}}_{L,k}\odot\mathbf{\bar{G}}_{L,k})\mathbf{z}_{L}^{\mathrm{H}}$, where $\mathbf{\bar{H}}_{L,k}\in\mathbb{C}^{N_{L}\times N_{L}}=\mathbf{H}_{L}\mathbf{w}_{k}\mathbf{w}_{k}^{\mathrm{H}}\mathbf{H}_{L}^{\mathrm{H}}$ and $\mathbf{\bar{G}}_{L,k}^{\mathrm{T}}\in\mathbb{C}^{N_{L}\times N_{L}}=\mathbf{g}_{L,k}^{\mathrm{H}}\mathbf{g}_{L,k}$, while $\mathbf{z}_{L}\in\mathbb{C}^{1\times N_{L}}$ is the reflection vector of $L$-RIS that composes non-zero diagonal elements of reflection matrix $\mathbf{\Omega}_{L}$. In a similar way, the denominator of $\eta_{k}$ can be rewritten in quadratic form as $\left\|\mathbf{g}_{L,k}\mathbf{\Omega}_{L}\mathbf{H}_{L}\mathbf{w}_{\hat{k}}\right\|_{\mathrm{F}}^{2}=\mathbf{z}_{L}(\mathbf{\bar{H}}_{L,k}\odot\mathbf{\bar{G}}_{L,\hat{k}})\mathbf{z}_{L}^{\mathrm{H}}$ and $\left\|\mathbf{g}_{M,k}\mathbf{\Omega}_{M}\mathbf{H}_{M}\mathbf{W}\right\|_{\mathrm{F}}^{2}=\mathbf{z}_{M}(\mathbf{\bar{H}}_{M,k}\odot\mathbf{\bar{G}}_{M,{k}})\mathbf{z}_{M}^{\mathrm{H}}$, where $\mathbf{\bar{G}}_{L,\hat{k}}^{\mathrm{T}}\in\mathbb{C}^{N_{L}\times N_{L}}=\mathbf{g}_{L,\hat{k}}^{\mathrm{H}}\mathbf{g}_{L,\hat{k}}$, $\mathbf{\bar{H}}_{M,k}\in\mathbb{C}^{N_{M}\times N_{M}}=\mathbf{H}_{M}\mathbf{R_{x}}\mathbf{H}_{M}^{\mathrm{H}}$ and $\mathbf{\bar{G}}_{M,k}^{\mathrm{T}}\in\mathbb{C}^{N_{M}\times N_{M}}=\mathbf{g}_{M,k}^{\mathrm{H}}\mathbf{g}_{M,k}$, while $\mathbf{z}_{M}\in\mathbb{C}^{1\times N_{M}}$ is the random interference vector of $M$-RIS that includes non-zero diagonal elements of malicious reflecting matrix $\mathbf{\Omega}_{M}$. Therefore, (27) can be rewritten as
$\eta_{k}=\frac{\operatorname{Tr}(\mathbf{C}_{L,k}\mathbf{Z}_{L})}{\sum_{\hat{k}}\operatorname{Tr}(\mathbf{C}_{L,\hat{k}}\mathbf{Z}_{L})+\operatorname{Tr}(\mathbf{C}_{M,\hat{k}}\mathbf{Z}_{M})+\sigma_{n}^{2}}$ (36)
@@
where $\mathbf{C}_{L,k}\in\mathbb{C}^{N_{L}\times N_{L}}=\mathbf{\bar{H}}_{L,k}\odot\mathbf{\bar{G}}_{L,k}$, $\mathbf{C}_{L,\hat{k}}=\mathbf{\bar{H}}_{L,k}\odot\mathbf{\bar{G}}_{L,\hat{k}}$ and $\mathbf{C}_{M,k}\in\mathbb{C}^{N_{M}\times N_{M}}=\mathbf{\bar{H}}_{M,k}\odot\mathbf{\bar{G}}_{M,k}$, while $\mathbf{Z}_{L}\in\mathbb{C}^{N_{L}\times N_{L}}=\mathbf{z}_{L}^{\mathrm{H}}\mathbf{z}_{L}$ and $\mathbf{Z}_{M}\in\mathbb{C}^{N_{M}\times N_{M}}=\mathbf{z}_{M}^{\mathrm{H}}\mathbf{z}_{M}$. Therefore, the legitimate phase optimization problem in (35) can be converted to the following quadratically constrained quadratic programming (QCQP) problem as follows
$\text{(P3)}\max_{\mathbf{Z}_{L}}\hskip 2.84544pt\min\eta_{k}$ (37a)
s.t.
$\operatorname{Tr}(\mathbf{C}_{L,k}\mathbf{Z}_{L})-\Big{(}\sum_{\hat{k}}\operatorname{Tr}(\mathbf{C}_{L,\hat{k}}\mathbf{Z}_{L})+\operatorname{Tr}(\mathbf{C}_{M,\hat{k}}\mathbf{Z}_{M})+\sigma_{n}^{2}\Big{)}\geq 0$
$\hskip 39.83368pt\mathbf{Z}_{L}\succeq 0$ (37c)
$\hskip 39.83368pt\mathbf{Z}_{L}(l,l)=1\hskip 5.69046pt\text{for}\hskip 5.69046ptl\in\{1,\cdots,N_{L}\}.$ (37d)

### v2

@@ II-C Communication Model
$\mathbf{H}_{n}=\sqrt{\frac{{L_{0}}}{d_{n}^{\alpha_{n}}}}\left({\sqrt{\frac{\kappa}{1+\kappa}}\mathbf{H}^{\text{LOS}}_{n}+\sqrt{\frac{1}{1+\kappa}}\mathbf{H}^{\text{NLOS}}_{n}}\right),$ (20)
$\mathbf{g}_{n,k}=\sqrt{\frac{{L_{0}}}{d_{n,k}^{\alpha_{k}}}}\left(\sqrt{\frac{\kappa}{1+\kappa}}\mathbf{g}_{n,k}^{\text{LOS}}+\sqrt{\frac{1}{1+\kappa}}\mathbf{g}_{n,k}^{\text{NLOS}}\right),$ (21)
@@
$\mathbf{H}^{\text{LOS}}_{n}=\mathbf{b}(\theta_{n}^{\text{RIS}},\phi_{n}^{\text{RIS}})^{\mathrm{H}}\mathbf{b}(\theta_{n}^{\text{BS}},\phi_{n}^{\text{BS}}),$ (22)
$\mathbf{g}^{\text{LOS}}_{n,k}=\mathbf{b}(\theta_{n,k}^{\text{RIS}},\phi_{n,k}^{\text{RIS}}),$ (23)
@@
$\mathbf{b}(\theta_{n}^{\text{BS}},\phi_{n}^{\text{BS}})=[1,\cdots,e^{j{\nu_{T}}(T_{x}-1)\cos(\theta^{\text{BS}}_{n})\cos(\phi^{\text{BS}}_{n})}],$ (24)
$\mathbf{b}(\theta_{n}^{\text{RIS}},\phi_{n}^{\text{RIS}})=\mathbf{b}^{x}(\theta_{n}^{\text{RIS}},\phi_{n}^{\text{RIS}})\otimes\mathbf{b}^{z}(\theta_{n}^{\text{RIS}},\phi_{n}^{\text{RIS}}),$ (25)
where $\mathbf{b}^{x}(\theta_{n}^{\text{RIS}},\phi_{n}^{\text{RIS}})\in\mathbb{C}^{1\times N_{n}^{x}}$ and $\mathbf{b}^{z}(\theta_{n}^{\text{RIS}},\phi_{n}^{\text{RIS}})\in\mathbb{C}^{1\times N_{n}^{z}}$ represent the steering vectors of the $n$-RIS along the $x$- and $z$-axes, respectively. $N_{x}^{n}$ and $N_{z}^{n}$ denote the number of reflecting elements along the $x$- and $z$-directions, and the total number of elements is given by $N_{n}=N_{x}^{n}\times N_{z}^{n}$ for $n\in\{L,M\}$. Therefore, $\mathbf{b}^{x}(\theta_{n}^{\text{RIS}},\phi_{n}^{\text{RIS}})$ and $\mathbf{b}^{z}(\theta_{n}^{\text{RIS}},\phi_{n}^{\text{RIS}})$ can be given as
$\mathbf{b}^{x}(\theta_{n}^{\text{RIS}},\phi_{n}^{\text{RIS}})=[1,\cdots,e^{j{\nu_{R}}(N_{n}^{x}-1)\cos(\theta^{\text{RIS}}_{n})\cos(\phi^{\text{RIS}}_{n})}],$ (26a)
$\mathbf{b}^{z}(\theta_{n}^{\text{RIS}},\phi_{n}^{\text{RIS}})=[1,\cdots,e^{j{\nu_{R}}(N_{n}^{z}-1)\sin(\phi^{\text{RIS}}_{n})}],$ (26b)
@@
$y_{k}(t)=\mathbf{g}_{L,k}\mathbf{\Omega}_{L}\mathbf{H}_{L}\mathbf{x}(t)+\mathbf{g}_{M,k}\mathbf{\Omega}_{M}\mathbf{H}_{M}\mathbf{x}(t)+n_{k}(t),$ (27)
@@
$\eta_{k}=\frac{\left|\mathbf{g}_{L,k}\mathbf{\Omega}_{L}\mathbf{H}_{L}\mathbf{w}_{k}\right|^{2}}{\sum_{\hat{k}}\left|\mathbf{g}_{L,k}\mathbf{\Omega}_{L}\mathbf{H}_{L}\mathbf{w}_{\hat{k}}\right|^{2}+\left\|\mathbf{g}_{M,k}\mathbf{\Omega}_{M}\mathbf{H}_{M}\mathbf{W}\right\|_{\mathrm{F}}^{2}+\sigma_{n}^{2}},$ (28)
@@ II-D Security Model
In the proposed scheme, the eavesdropper UAV target performs two separate functions: it (i) attempts to intercept the communication signals through a direct BS and UAV propagation channel, and (ii) carries a malicious RIS whose role is to degrade the reception at legitimate users through an interference attack. Accordingly, using (8), the eavesdropping link is modeled as the direct link between the BS eavesdropper channel and the received signal at the eavesdropper UAV target at the $t$-th time index can be expressed as
$y_{E}(t)=\sqrt{\beta_{E}}\mathbf{a}(\theta_{E}^{\text{BS}},\phi_{E}^{\text{BS}})\mathbf{x}(t)+n_{E}(t),$ (29)
@@
$\eta_{E,k}=\frac{\left|\mathbf{a}(\theta_{E}^{\text{BS}},\phi_{E}^{\text{BS}})\mathbf{w}_{k}\right|^{2}}{\sum_{\hat{k}}\left|\mathbf{a}(\theta_{E}^{\text{BS}},\phi_{E}^{\text{BS}})\mathbf{w}_{\hat{k}}\right|^{2}+\sigma_{n}^{2}},$ (30)
@@
$S_{R,k}=[R_{k}-R_{E,k}]^{+},$ (31)
where $x]^{+}=\max(x,0)$. $R_{k}=\log_{2}(1+\eta_{k})$ is the achievable rate of user Uk and $R_{E,k}=\log_{2}(1+\eta_{E,k})$ represents the eavesdropping rate at for Uk.
@@ III Problem Formulation and Proposed Solution
It is crucial to emphasize that in the proposed scheme, it is assumed that the BS is unaware of the malicious RIS mounted on the $E$-UAV whose reflection matrix $\mathbf{\Omega}_{M}$ in (2) using randomly generated phase shifts. Therefore, the received signal at $U_{k}$ that perceived by the BS differs from (27) and can be given as [2]
$\hat{y}_{k}(t)=\mathbf{g}_{L,k}\mathbf{\Omega}_{L}\mathbf{H}_{L}\mathbf{x}(t)+n_{k}(t).$ (32)
Therefore, using (32), the BS formulates the instantaneous SINR of the Uk as
$\hat{\eta}_{k}=\frac{\left|\mathbf{g}_{L,k}\mathbf{\Omega}_{L}\mathbf{H}_{L}\mathbf{w}_{k}\right|^{2}}{\sum_{\hat{k}}\left|\mathbf{g}_{L,k}\mathbf{\Omega}_{L}\mathbf{H}_{L}\mathbf{w}_{\hat{k}}\right|^{2}+\sigma_{n}^{2}}.$ (33)
Accordingly, using (33), from the BS perspective, the secrecy rate of the $k$-th legitimate user is calculated as
$\hat{S}_{R,k}=[\hat{R}_{k}-R_{E,k}]^{+}$ (34)
where $\hat{R}_{k}=\log_{2}(1+\hat{\eta}_{k})$ is assumed to be the instantaneous achievable rate of Uk.
Therefore, based on the above calculations at the BS, the reflection coefficients of the legitimate RIS and the BS transmit beamforming matrix are jointly optimized to maximize the secrecy rate of the overall system. To address this, the following optimization problem is formulated:
$\text{(P1)}\hskip 11.38092pt\max_{\mathbf{W},\mathbf{\Omega}_{L}}\hskip 5.69046pt{\hat{S}_{R}}$ (35a)
$\hskip 42.67912pt\text{s.t.}\hskip 5.69046pt\operatorname{Tr}(\mathbf{WW}^{\mathrm{H}})\leq P_{T},$ (35b)
$\hskip 65.44142pt\gamma_{i}\leq\rho_{s}\hskip 8.5359pt\text{for}\hskip 5.69046pti\in\{L,E\},$ (35c)
$\hskip 65.44142pt|e^{j\varphi_{l}}|=1,\hskip 8.5359pt\text{for}\hskip 5.69046ptl\in\{1,\cdots,N_{L}\},$ (35d)
where $\hat{S}_{R}=\sum\limits_{k=1}^{K}\hat{S}_{R,k}$} is the achievable sum secrecy rate of legitimate users and $\rho_{s}$ is the maximum achievable sensing SINR rate. Since both legitimate and eavesdropper UAVs communicate directly with the BS without any reflected signals from RISs, the beamforming optimization can simply be independent of $L$-RIS configuration. This allows the problem (P1) to be decomposed into a two-stage optimization framework that first optimizing the transmit beamforming matrix $\mathbf{W}$, and later optimizing the $L$-RIS reflection matrix $\mathbf{\Omega}_{L}$.
Accordingly, the joint design problem in (P1) involves the optimization of both the BS transmit beamforming matrix $\mathbf{W}$ and the $L$-RIS reflection matrix $\mathbf{\Omega}_{L}$. Because the direct BS-UAV links are not influenced by RIS reflections, the two variables become separable, enabling a sequential optimization framework. In the first stage, (P2) is formulated to optimize only the beamforming matrix $\mathbf{W}$ by maximizing the minimum sensing SINR of the two UAV targets under the transmit power and sensing quality constraints. Once the optimum beamforming matrix $\mathbf{W}$ is obtained, it yields a second-stage optimization denoted as (P3) that focuses solely on determining the constructive RIS matrix $\mathbf{\Omega}_{L}$ so as to maximize the achievable sum secrecy rate. This two-step decomposition cooperatively contributes to the overall goal of (P1).
@@ III-A Beamforming Optimization
$\text{(P2)}\hskip 11.38092pt\max_{\mathbf{W}}{\min_{i}}\hskip 5.69046pt\gamma_{i}$ (36a)
$\hskip 56.9055pt{\text{s.t.}}\hskip 5.69046pt\gamma_{i}\leq\rho_{s},$ (36b)
$\hskip 76.82234pt\operatorname{Tr}(\mathbf{WW}^{\mathrm{H}})\leq P_{T}.$ (36c)
In (P2), the sensing SINR of the $i$-UAV target is maximized while ensuring that the minimum sensing SINR remains above the acceptable threshold $\rho_{s}$, thereby guaranteeing reliable sensing performance.
For $\mathbf{A}_{i}\in\mathbb{C}^{T_{x}\times T_{x}}=\beta_{i}\mathbf{a}(\theta^{\text{BS}}_{i},\phi^{\text{BS}}_{i})^{\mathrm{H}}\mathbf{a}(\theta^{\text{BS}}_{i},\phi^{\text{BS}}_{i})$ being the round trip channel between BS and $i$-UAV target, then sensing SINR $\gamma_{i}$ in (11) can be re-expressed in the following quadratic form
$\gamma_{i}=\frac{\operatorname{Tr}(\mathbf{A}_{i}^{\mathrm{H}}\mathbf{A}_{i}\mathbf{R_{x}})}{\operatorname{Tr}(\mathbf{A}_{j}^{\mathrm{H}}\mathbf{A}_{j}\mathbf{R_{x}})+\sigma_{n}^{2}},$ (37)
where $j\neq i\in\{L,E\}$. Therefore, using a SDR-based approach, (P2) can be reformulated as
$\text{(P2)}\hskip 11.38092pt\max_{\mathbf{R_{x}}}\hskip 2.84544pt{\min_{i}}\hskip 5.69046pt\gamma_{i}$ (38a)
$\hskip 54.06006pt{\text{s.t.}}\hskip 8.5359pt\gamma_{i}\leq\rho_{s},$ (38b)
$\hskip 73.97733pt\operatorname{Tr}(\mathbf{A}_{L}^{\mathrm{H}}\mathbf{A}_{L}\mathbf{R_{x}})\geq\operatorname{Tr}(\mathbf{A}_{E}^{\mathrm{H}}\mathbf{A}_{E}\mathbf{R_{x}})+\sigma_{n}^{2},$ (38c)
$\hskip 73.97733pt\operatorname{Tr}(\mathbf{R_{x}})\leq P_{T}.$ (38d)
It is important to note that, as reflected in constraint (38c), the BS prioritizes the $L$-UAV target while maintaining a minimum sensing SINR threshold $\rho_{s}$ for the $E$-UAV. After that, using the SeDuMi solver within the CVX optimization toolbox [8], (P2) can be effectively solved. Subsequently, eigenvalue decomposition (EVD) is applied to recover the beamforming matrix $\mathbf{W}$ [34].
Moreover, in order to evaluate the spatial selectivity of the optimized beamforming matrix, ensuring that the transmitted energy is effectively concentrated toward the desired UAV target, $L$-UAV, while minimizing interference in other directions, the beampattern gain of the $i$-th target in direction of $\{\theta^{\text{BS}}_{i},\phi^{\text{BS}}_{i}\}$ angles can be calculated using steering vectors $\mathbf{a}(\theta^{\text{BS}}_{i},\phi^{\text{BS}}_{i})$ in (9) as
$\mathcal{P}_{i}=\mathbf{a}(\theta^{\text{BS}}_{i},\phi^{\text{BS}}_{i})^{\mathrm{H}}\mathbf{R_{x}}\mathbf{a}(\theta^{\text{BS}}_{i},\phi^{\text{BS}}_{i}).$ (39)
@@ III-B Legitimate RIS Phase Optimization
In this subsection, after the beamforming matrix $\mathbf{W}$ is obtained, the BS optimizes the reflection coefficients of $L$-RIS to maximize the secrecy rate of the $K$ legitimate users. At the BS, the secrecy rate $\hat{S}_{R}$ in (34) is computed as the total sum of achievable secrecy rates across all legitimate users. Since eavesdropper SINR of $E$-UAV target (30) across users are independent from $L$-RIS reflection, the $\boldsymbol{\Omega}_{L}$ (1) can be optimized to enhance $\hat{S}_{R}$ by maximizing sum of achievable rates of all users $\sum_{k}\hat{R}_{k}$. Therefore, considering (33), the optimization of $\mathbf{\Omega}_{L}$ in (P1) can be converted to following problem:
$\text{(P3)}\hskip 11.38092pt\max_{\mathbf{\Omega}_{L}}\hskip 2.84544pt{\min_{k}}\hskip 2.84544pt{\hat{\eta}_{k}}$ (40a)
@@
As a first step, the instantaneous communication SINR given in (33) is re-expressed in a quadratic form by applying appropriate mathematical modifications. Applying trace equality [34], the numerator in (33) can be rewritten as $\left|\mathbf{g}_{L,k}\mathbf{\Omega}_{L}\mathbf{H}_{L}\mathbf{w}_{k}\right|^{2}=\operatorname{Tr}(\mathbf{\bar{H}}_{L,k}\mathbf{\Omega}_{L}^{\mathrm{H}}\mathbf{\bar{G}}_{L,k}^{\mathrm{T}}\mathbf{\Omega}_{L})=\mathbf{z}_{L}(\mathbf{\bar{H}}_{L,k}\odot\mathbf{\bar{G}}_{L,k})\mathbf{z}_{L}^{\mathrm{H}}$, where $\mathbf{\bar{H}}_{L,k}\in\mathbb{C}^{N_{L}\times N_{L}}=\mathbf{H}_{L}\mathbf{w}_{k}\mathbf{w}_{k}^{\mathrm{H}}\mathbf{H}_{L}^{\mathrm{H}}$ and $\mathbf{\bar{G}}_{L,k}^{\mathrm{T}}\in\mathbb{C}^{N_{L}\times N_{L}}=\mathbf{g}_{L,k}^{\mathrm{H}}\mathbf{g}_{L,k}$, while $\mathbf{z}_{L}\in\mathbb{C}^{1\times N_{L}}$ is the reflection vector of $L$-RIS that composes non-zero diagonal elements of reflection matrix $\mathbf{\Omega}_{L}$. In a similar way, the denominator of $\hat{\eta}_{k}$ can be rewritten in quadratic form as $\left|\mathbf{g}_{L,k}\mathbf{\Omega}_{L}\mathbf{H}_{L}\mathbf{w}_{\hat{k}}\right|^{2}=\mathbf{z}_{L}(\mathbf{\bar{H}}_{L,k}\odot\mathbf{\bar{G}}_{L,\hat{k}})\mathbf{z}_{L}^{\mathrm{H}}$ where $\mathbf{\bar{G}}_{L,\hat{k}}^{\mathrm{T}}\in\mathbb{C}^{N_{L}\times N_{L}}=\mathbf{g}_{L,\hat{k}}^{\mathrm{H}}\mathbf{g}_{L,\hat{k}}$, Therefore, (33) can be rewritten as
$\hat{\eta}_{k}=\frac{\operatorname{Tr}(\mathbf{C}_{L,k}\mathbf{Z}_{L})}{\sum_{\hat{k}}\operatorname{Tr}(\mathbf{C}_{L,\hat{k}}\mathbf{Z}_{L})+\sigma_{n}^{2}},$ (41)
@@
where $\mathbf{C}_{L,k}\in\mathbb{C}^{N_{L}\times N_{L}}=\mathbf{\bar{H}}_{L,k}\odot\mathbf{\bar{G}}_{L,k}$, $\mathbf{C}_{L,\hat{k}}=\mathbf{\bar{H}}_{L,k}\odot\mathbf{\bar{G}}_{L,\hat{k}}$, while $\mathbf{Z}_{L}\in\mathbb{C}^{N_{L}\times N_{L}}=\mathbf{z}_{L}^{\mathrm{H}}\mathbf{z}_{L}$ Therefore, the legitimate phase optimization problem in (40) can be converted to the following quadratically constrained quadratic programming (QCQP) problem as follows
$\text{(P3)}\hskip 8.5359pt\max_{\mathbf{Z}_{L}}\hskip 5.69046pt{\min_{k}}\hskip 2.84544pt\hat{\eta}_{k}$ (42a)
$\hskip 31.2982pt\text{s.t.}\hskip 5.69046pt\operatorname{Tr}(\mathbf{C}_{L,k}\mathbf{Z}_{L})-\Big(\sum_{\hat{k}}\operatorname{Tr}(\mathbf{C}_{L,\hat{k}}\mathbf{Z}_{L})+\sigma_{n}^{2}\Big)\geq 0,$ (42b)
$\hskip 51.21504pt\mathbf{Z}_{L}\succeq 0,$ (42c)
$\hskip 51.21504pt\mathbf{Z}_{L}(l,l)=1\hskip 5.69046pt\text{for}\hskip 5.69046ptl\in\{1,\cdots,N_{L}\}.$ (42d)